OpenStudy (jennychan12):

What are the equations of the horizontal asymptotes of y = (4e^x+7)/(e^x-1) in the xy plane?

OpenStudy (jennychan12):

\[y = \frac{ 4e^x + 7}{ e^x - 1 }\]

OpenStudy (jennychan12):

I know one horizontal asymptote is y = 4. But there are two... The other one is y = -7, but how do you get that one? I know you can break up the numerator so that \[y = \frac{ 4e^x }{ e^x - 1 } + \frac{ 7 }{ e^x - 1 }\] but how would you get the -7?

OpenStudy (anonymous):

as \(x\to -\infty\) you have \(e^x\to 0\)

OpenStudy (jennychan12):

ohhhhh i see it now. -_- wow. okay, thanks.

OpenStudy (anonymous):

so \[\lim_{x\to -\infty}\frac{ 4e^x + 7}{ e^x - 1 }=\frac{7}{-1}\]

OpenStudy (anonymous):

yw